Calculate the number of electron which will together weigh 1 gram .

Grade:11

1 Answers

Vikas TU
4 years ago
Mass of one electron = 9.10939×10^-31 kg
Number of electrons that weight 9.10939×10^-31 kg= 1
Number of electrons that weight 1 gram=(1×10^-3 kg)
=0.1098×10^-3+31
=0.1098×10^28
=1.098×10^27
Hope this could be the answer.